# Stellar-gas kinematic misalignments in eagle: lifetimes and longevity of misaligned galaxies

### v1

@@ 4 Sample
    A maximum spatial separation between the stellar and gasSF centres of mass within $<2$ pkpc. This allows us to exclude galaxies featuring highly perturbed cold gas reservoirs that are only loosely associated with the galaxy (see Appendix D in Trayford & Schaye, 2019, for examples of this in high-mass galaxies).
@@
    A maximum uncertainty of $<30^{\circ}$ for the misalignment angle using boostrap resampling. This criterion is motivated by the presence of systems that are heavily dispersion-dominated and/or feature relatively few gasSF particles (as is typical for many ETGs).
Unlike Casanueva et al. (2022), we do not impose a direct limit on dispersion-dominated systems through the use of the stellar spin parameters (see Emsellem et al., 2011). Instead, we use bootstrap resampling and take the $1\sigma$ ($68\%$) percentiles to estimate the uncertainty in the misalignment angle. This uncertainty method will naturally scale with the degree of stellar dispersion in the galaxy, but will also remove a handful of galaxies with chaotic gasSF kinematics that are not captured by the stellar spin parameter. Our uncertainty limit of $<30^{\circ}$ is motivated by similar uncertainty thresholds used in observational studies (e.g. Davis et al., 2011). At $z=0.1$, this criterion reduces our sample size by $\approx 6.5\%$. The majority of the excluded objects are dispersion-dominated ETGs with low gasSF particle counts.
At $z=0.1$, this gives us a sample of 5428 galaxies. Similar numbers are obtained at redshifts up to $z\approx 1$. Overall, the adopted criteria produce galaxy samples that do not show significant bias towards any galaxy mass, for a range of snapshots. For instance, Figure 1 shows the stellar mass function of our final sample at $z=0.1$, alongside the total galaxy population in the Ref-L100N1504 reference simulation, which shows that the two populations have a similar mass distribution down to our sample limit.
@@ 5.1.1 Star-forming gas and star-forming gas fraction
$n^{*}_{\rm{H}}=10^{-1}\rm{cm}^{-3}\left(\frac{Z}{0.002}\right)^{-0.64},$ (3)
where $Z$ is the metal mass fraction (i.e. the gas metallicity). This metallicity-dependant threshold creates a lower threshold for star formation in regions of higher metallicity, as expected by the increased cooling efficiency and shielding provided by dust in enriched gas in the ISM. The exact SFR of gas particles is handled by the subgrid model and described in detail in Schaye & Dalla Vecchia (2008).
We define the star-forming gas fraction of a galaxy as {ceqn}
@@ 5.1.2 Morphology
We make use of the ratio of the co-rotational kinetic energy fraction, $\kappa_{\rm{co}}$, to distinguish between galaxy morphologies. This describes the ratio of total kinetic energy ($K$) that is invested in co-rotation ($K^{\rm{rot}}_{\rm{co}}$) and is given by {ceqn}
@@
where the sum is over all particles of a given type within a sphere of radius 30 pkpc centred on the minimum potential, $m_{i}$ is the particle mass, $L_{\rm{z},\textit{i}}$ is the net angular momentum, and $R_{i}$ is the radius from the centre of potential in the plane normal to the rotation axis of the given particle type (Correa et al., 2017). We denote stellar co-rotational kinetic energy fraction as $\kappa_{\rm{co}}^{*}$. Similarly, we define the gasSF co-rotational kinetic energy fraction as $\kappa_{\rm{co}}^{\rm{SF}}$ using all gasSF particles within a sphere of radius $r_{50}$.
Correa et al. (2017) found that ‘blue cloud’ discy star-forming galaxies and ‘red sequence’ quiescent spheroidal galaxies can be distinguished using a threshold of $\kappa_{\rm{co}}^{*}\approx 0.4$ for stellar particles in eagle. We use this as a proxy for morphology to distinguish between LTGs with $\kappa_{\rm{co}}^{*}>0.4$, and ETGs with $\kappa_{\rm{co}}^{*}<0.4$. Similarly to Jiménez et al. (2023), we find by visual inspection that thin gas discs can be broadly described by $\kappa_{\rm{co}}^{\rm{SF}}\gtrsim 0.7$.
To characterise the torquing timescale, we make use of a further proxy of galaxy morphology, the stellar ellipticity, which we define as {ceqn}
@@
where $a$ and $c$ are the major and minor axes of the ellipse, respectively. As opposed to $\kappa_{\rm{co}}$, these use the square root of the eigenvalues of the mass distribution tensor to characterise the morphology of a galaxy within a sphere of 30 pkpc centred on the potential minimum (for details see Correa et al., 2017).
@@ 5.1.3 Misalignment angles
We assume the kinematics of the particles are captured by the net angular momentum vector. For galaxies meeting our selection criteria, the angular momentum of a component is given by {ceqn}
@@
where the sum is over all particles of a specified type within $r_{50}$, $m_{i}$ is the particle mass, $\mathbf{r}_{i}$ and $\mathbf{v}_{i}$ are the physical position and velocity of the particle, and $\mathbf{r}_{\rm{COM}}$ and $\mathbf{v}_{\rm{COM}}$ are the position and peculiar velocity of the stellar centre of mass within $r_{50}$. For dark matter (DM) particles, we extract kinematic properties within a spherical aperture of 30 pkpc.
@@
We assume the most widely used classification to interpret galaxy kinematics (e.g. Davis et al., 2011), classifying as aligned those galaxies with stellar-gas kinematic angles $<30^{\circ}$, misaligned those with angles $>30^{\circ}$. Among the misaligned galaxy population, we make a further distinction between unstable misalignments with angles $[30^{\circ}-150^{\circ}]$ and stable misalignments (i.e. counter-rotation) with angles $>150^{\circ}$.
@@ 5.2 External properties
We assume that the environment of a galaxy is described by the halo mass, $M_{\rm{200}}$, bound within a radius 200 times the critical density as identified in subfind.
We define galaxy properties related to mergers using the merger trees by Qu et al. (2017). Merger ratios are calculated from values in the merger tree at the point of the maximum stellar mass of the progenitor galaxy within the past 2 Gyr to account for stellar stripping prior to infall (see Rodriguez-Gomez et al., 2015; Qu et al., 2017). We define the mass ratio of a merger as {ceqn}
@@
where ${M^{S}_{*}}$ and $M^{P}_{*}$ are the stellar masses of the secondary (less massive) and primary (most massive) galaxies, respectively. Major and minor mergers are defined with mass ratios of $\mu_{\rm{*}}\geq 0.3$ and $0.3\geq\mu_{\rm{*}}\geq 0.1$ respectively. We assume that ratios of $\mu_{\rm{*}}<0.1$ are direct accretion events (sometimes called micro-mergers).
@@ 5.3.1 Relaxation times
In order to estimate the relaxation time of an unstable misalignment we require galaxies to meet our sample criteria over consecutive snapshots from misalignment formation to relaxation. We also exclude unstable misalignments for which we cannot establish a formation and relaxation point within $10<z<$ 1. By limiting ourselves to galaxies with a complete ‘relaxation window’, we extract a smaller sample of relaxations but are able to establish a more reliable estimate of relaxation times. Additionally, our redshift range naturally places an upper limit on the relaxation time that we can accurately measure in our sample, with $t_{\mathrm{relax}}\approx 40t_{\mathrm{dyn}}$ approaching the $\approx 8$ Gyr window between $10<z<$ 1, for a typical ETG.
@@
For this work, we adopt a similar definition and approach as van de Voort et al. (2015), modified for use on a more diverse population of galaxies and misalignments. Our aim here is to generalise relaxation timescales of unstable misalignments for comparison with theoretical predictions (see Section 2), complementing the approach of Khim et al. (2021). As an unstable misalignment may be formed from an existing long-lived counter-rotating system, we define a separate kinematically-stable regime of $\psi_{\rm{3D}}<20^{\circ}$ and $\psi_{\rm{3D}}>160^{\circ}$. This is motivated by visual inspection of typical galaxies that are found to mostly reside within this regime, with only few outliers caused by random fluctuations. We then require any galaxy leaving the kinematically-stable regime to enter the unstable misaligned regime ($30^{\circ}-150^{\circ}$) in order to be classified as significant deviation from the kinematically-stable regime [^3] Following our existing method, we note that changing the stable regime limits to $\psi_{\rm{3D}}<30^{\circ}$ and $\psi_{\rm{3D}}>150^{\circ}$, and the unstable misaligned regime to $[45^{\circ}-135^{\circ}]$ does not meaningfully change our results.. In most cases, the first snipshot in the unstable regime also corresponds to the first snipshot in the unstable misaligned regime. We define the beginning of a relaxation from the first snipshot in the kinematically unstable regime.
As some misalignments may experience a small degree of scatter during relaxation to the stable regime, we introduce a new parameter, the latency time, $t_{\rm{lat}}$. This is defined as the minimum amount of time a galaxy must have spent in the stable regime for the unstable misalignment to have truly settled. We set this value to $t_{\rm{lat}}=0.1$ Gyr ($\gtrsim 1$ snipshots), again from visual inspection. We define the end of a relaxation as the first snipshot upon returning to the stable regime, so long as the galaxy remains in the stable regime for period of at least another $t_{\rm{lat}}$. The relaxation time, $t_{\rm{relax}}$, is then defined as the time between the snipshots corresponding to the beginning and end of the misalignment. This gives us an error of approximately ±$\pm 0.12$ Gyr on $t_{\rm{relax}}$.
By applying this criterion, we aim to find a balance between extracting numerous short-duration relaxations that may be associated with small fluctuations from a larger overall relaxation, and extracting clearly separate unstable misalignment events as one long relaxation. We demonstrate the reliability of this criterion by showing two example galaxies in Section 6.2 and discuss its implications on our relaxation timescale distribution in Appendix A.
@@ 5.3.2 Dynamical and torquing times
The $t_{\rm{torque}}$ of a galaxy can vary between snipshots due to changes in the gas disc radius and stellar ellipticity. The gas disc radius and stellar ellipticity were found to vary within a standard deviation of $\sim 10\%$ between misalignment formation and relaxation in our sample. We express the relaxation time in terms of time passed in Gyr and as normalised values using the the time-averaged dynamical and torquing time taken over the duration of the misalignment, denoted as $\bar{t}_{\rm{dyn}}$ and $\bar{t}_{\rm{torque}}$ , respectively. Following theoretical predictions (see Section 2), normalised relaxation timescales above $t_{\rm{relax}}/\bar{t}_{\rm{torque}}\gtrsim 1$ would be indicative of processes acting to preserve the duration of the misalignment. To account for approximations made, we define long-duration unstable misalignments as those with $t_{\rm{relax}}/\bar{t}_{\rm{torque}}>3$.
@@ 6.1 Misalignment distributions
For a brief comparison of distribution angles with observations, we make use of projected misalignment angles as viewed from the $z$-axis ($\psi_{\rm{2D}}$), within a projected half-mass radii ($r_{50,z}$) in the $x-y$ plane, though our choice of axis is arbitrary as galaxies within eagle have no preferential orientation.
@@
Figure 2 shows the projected and 3D misalignment distributions for LTGs (top) and ETGs (bottom) at $z=0.1$, with uncertainties given as Poisson errors. For values in projection, we find $38.9\pm 5.3\%$ of ETGs and $7.9\pm 1.7\%$ of LTGs display stellar-gas kinematic misalignments. Both the fractions of misalignments and the general shape of the misalignment distributions are in good agreement with recent observational results by Bryant et al. (2019). These were obtained using the SAMI galaxy survey (Croom et al., 2012; Bryant et al., 2015), with observational misalignment fractions of $\approx 33\%$ (ETG) and $\approx 5\%$ (LTG), but with a stellar mass range extending a larger range of galaxy masses $10^{8}\lesssim M_{*}\lesssim 10^{12}$ M. Our results are also in good agreement with a number of other observational studies (e.g. Davis et al., 2011; Barrera-Ballesteros et al., 2014, 2015; Chen et al., 2016; Jin et al., 2016; Ristea et al., 2022; Raimundo et al., 2023), as well as with predictions from cosmological simulations (e.g. Khim et al., 2020; Casanueva et al., 2022).
@@
We find a small peak in the counter-rotating region for both morphologies, with projected counter-rotating fractions of $10.9\pm 1.3\%$ and $3.3\pm 0.5\%$ for ETGs and LTGs, respectively. These were not reported in previous results in eagle (see Casanueva et al., 2022), or in the horizon-agn simulation (see Khim et al., 2020). We visually inspected our sample of $\psi_{\rm{2D}}>170^{\circ}$ misalignments and confirmed these systems to be physical and misaligned in both 3D and projection. As a whole, we find counter-rotation present in $5.7\pm 0.6\%$ of our sample. To test the significance of this result, we perform a Kolmogorov–Smirnov (KS) test in the range $[135^{\circ}-180^{\circ}]$ and compare this to a uniform distribution in the same range. We obtain KS-statistic $=0.19$ and p-value $=1.8\times 10^{-9}$ for ETGs and KS-statistic $=0.41$ and p-value $=3.3\times 10^{-18}$ for LTGs, indicating a statistically significant population of counter-rotators.
@@
Figure 2: Projected 2D (cyan/orange) and 3D (blue/red) stellar-gasSF misalignment angle distributions for LTGs (top) and ETGs (bottom) for our sample of misalignments at $z\approx 0.1$. The vertical dashed black line denotes the boundary between aligned ($<30^{\circ}$) and misaligned galaxy populations ($>30^{\circ}$). Observational results from Bryant et al. (2019) are included in grey and should be compared against $\psi_{\rm{2D}}$ distributions. Errors are given as Poisson uncertainties. The simulated and observed galaxies exhibit fairly similar distributions.
@@
We find a 3D misalignment fraction of $44.9\pm 5.4\%$ for ETGs and $7.2\pm 0.5\%$ for LTGs. The overall shape of the $\psi_{\rm{3D}}$ misalignment distribution is similar to that of projected values, being almost identical for LTGs but with a smaller fraction of aligned ETGs in the range $[0^{\circ}-10^{\circ}]$. We also see a small decrease in the fraction of counter-rotating systems when measured in 3D, but the number of counter-rotating systems remains significant with respect to to a uniform distribution using a KS test. These results suggest that although bias from projected viewing angles can inflate the number of aligned ETGs, the use of projected angles does not affect the overall shape or significantly reduce the number of misaligned systems.
@@
To justify our decision to consider both the co-rotating and counter-rotating regimes as stable, we show in Figure 3 how the fraction of aligned, misaligned, and counter-rotating systems varies for all snipshots between $10<z<$ 1. We consider the morphology independently and percentages are given with respect to the total number of galaxies of that morphological sub-sample.
For both ETGs and LTGs we see a slight upward trend in the misalignment fraction toward $z=0$. This increase appears to result overwhelmingly from the build-up of long-lived counter-rotating systems. This is seen most clearly for ETGs, in which the $\approx 5.1\%$ increase in misalignment fraction is captured by an $\approx 5.7\%$ increase in counter-rotating fraction. The build-up of these systems appears to indicate that counter-rotating systems are in fact long-lasting and dynamically stable, in agreement with theoretical predictions (Tohline et al., 1982; Lake & Norman, 1983). We thus decide to treat these as a separate stable regime among the misaligned galaxy population. Some fraction of these systems may return to a co-rotating state upon another stochastic gas accretion event or undergo an apparent stellar angular momentum flip as the gasSF forms a counter-rotating stellar disc that becomes dominant. For this latter process to occur, however, the masses of the counter-rotating gas and the stars have to be similar and the system must remain counter-rotating for long enough to form a dominant population of counter-rotating stars (for an observational counterpart see e.g. NGC 4550; Rubin et al., 1992; Rix et al., 1992; Wiklind & Henkel, 2001). We note that we see these apparent stellar angular momentum flips for several of our galaxies, as shown and discussed in Section 6.2.
@@
Figure 3: Stellar-gasSF misalignment angle fractions for LTGs (solid lines) and ETGs (dash-dotted lines) between $10<z<$ 1. Percentages are given with respect to the total galaxies of a given morphology at a given redshift. We classify galaxies as aligned with $\psi_{\rm{3D}}<30^{\circ}$ (blue), misaligned galaxies with $\psi_{\rm{3D}}>30^{\circ}$ (red), and counter-rotating galaxies with $\psi_{\rm{3D}}>150^{\circ}$ (purple). Errors are given as Poisson uncertainties (shaded regions). We see a slight upward trend in the misalignment fraction for both ETGs and LTGs across this redshift range, coinciding with an increase in counter-rotating fraction for both galaxy populations.
The longevity of counter-rotating systems in eagle is also in agreement with results obtained using other cosmological hydrodynamical simulations, such as those reported by Starkenburg et al. (2019). They used the IllustrisTNG (Nelson et al., 2019) hydrodynamical simulations to investigate the longevity of counter-rotating systems, finding these to be long-lasting relative to other orientations. We note that carrying out a detailed analysis of this behaviour is beyond the scope of this work.

### v2

@@ 4 Sample
    A maximum spatial separation between the stellar and gasSF centres of mass within <2<2 pkpc. This allows us to exclude galaxies featuring highly perturbed cold gas reservoirs that are only loosely associated with the galaxy (see Appendix D in Trayford & Schaye, 2019, for examples of this in high-mass galaxies).
@@
    A maximum uncertainty of $30^{\circ}$ for the misalignment angle using boostrap resampling. This criterion is motivated by the presence of systems that are heavily dispersion-dominated and/or feature relatively few gasSF particles (as is typical for many ETGs). Changing this limit to $10^{\circ}$ leaves our results and conclusions unchanged.
Unlike Casanueva et al. (2022), we do not impose a direct limit on dispersion-dominated systems through the use of the stellar spin parameters (see Emsellem et al., 2011). Instead, we use bootstrap resampling and take the $1\sigma$ ($68$ per cent) percentiles to estimate the uncertainty in the misalignment angle. This uncertainty method will naturally scale with the degree of stellar dispersion in the galaxy, but will also remove a handful of galaxies with chaotic gasSF kinematics that are not captured by the stellar spin parameter. Our uncertainty limit of $30^{\circ}$ is motivated by similar uncertainty thresholds used in observational studies (e.g. Davis et al., 2011). At $z=0.1$, this criterion reduces our sample size by $\approx 6.5$ per cent. The majority of the excluded objects are dispersion-dominated ETGs with low gasSF particle counts.
At $z=0.1$, this gives us a sample of 5428 galaxies. Similar numbers are obtained at redshifts up to $z\approx 1$. Overall, the adopted criteria produce galaxy samples that do not show significant bias towards any galaxy mass, for a range of snapshots. For instance, Figure 1 shows the stellar mass function of our final sample at $z=0.1$, alongside the total galaxy population in the Ref-L100N1504 reference simulation, which shows that the two populations have a similar mass distribution down to our sample limit. Additionally, this sample shows no bias in $M_{*}$-SFR (among galaxies with SFR >0>0) and no bias in halo mass (not shown).
@@ 5.1.1 Star-forming gas and star-forming gas fraction
We define the star-forming gas fraction of a galaxy as \@mathmargin
@@ 5.1.2 Morphology
eagle successfully reproduces the observed bimodal galaxy population of ‘red sequence’ spheroidal quiescent galaxies (ETGs) and ‘blue cloud’ discy star-forming galaxies (LTGs; e.g. Trayford et al., 2015; Correa et al., 2017; Thob et al., 2019; Pfeffer et al., 2023).
Correa et al. (2017) found that these populations can be effectively distinguished using the ratio of the co-rotational kinetic energy fraction, $\kappa_{\rm{co}}$, as a proxy for morphology. This describes the ratio of total kinetic energy ($K$) that is invested in co-rotation ($K^{\rm{rot}}_{\rm{co}}$) and is given by \@mathmargin
@@
where the sum is over all particles of a given type within a sphere of radius 30 pkpc centred on the minimum potential, $m_{i}$ is the particle mass, $L_{\rm{z},\textit{i}}$ is the net angular momentum, and $R_{i}$ is the radius from the centre of potential in the plane normal to the rotation axis of the given particle type (Correa et al., 2017). We denote stellar co-rotational kinetic energy fraction as $\kappa_{\rm{co}}^{*}$. Following Correa et al. (2017) and Thob et al. (2019), we use $\kappa_{\rm{co}}^{*}$ as a proxy for morphology to distinguish between ETGs ($\kappa_{\rm{co}}^{*}<0.4$) and LTGs ($\kappa_{\rm{co}}^{*}>0.4$).
Similarly, we define the gasSF co-rotational kinetic energy fraction as $\kappa_{\rm{co}}^{\rm{SF}}$ using all gasSF particles within a sphere of radius $r_{50}$. We find by visual inspection that thin gas discs can be broadly described by $\kappa_{\rm{co}}^{\rm{SF}}\gtrsim 0.7$, which is consistent to existing literature (e.g. Jiménez et al., 2023).
To characterise the torquing timescale, we make use of a further proxy of galaxy morphology, the stellar ellipticity, using the iterative scheme outlined in Thob et al. (2019). We define the stellar ellipticity as \@mathmargin
@@
where $a$ and $c$ are the major and minor axes of the ellipse, respectively. As opposed to $\kappa_{\rm{co}}$, these use the square root of the eigenvalues of the stellar mass distribution tensor to characterise the morphology of a galaxy within an initial spherical aperture of 30 pkpc (for details see Thob et al., 2019).
@@ 5.1.3 Misalignment angles
We assume the kinematics of the particles are captured by the net angular momentum vector. For galaxies meeting our selection criteria, the angular momentum of a component is given by \@mathmargin
@@
where the sum is over all particles of a specified type within $r_{50}$, $m_{i}$ is the particle mass, $\mathbf{r}_{i}$ and $\mathbf{v}_{i}$ are the physical position and velocity of the particle, and $\mathbf{r}_{\rm{COM}}$ and $\mathbf{v}_{\rm{COM}}$ are the position and peculiar velocity of the stellar centre of mass within $r_{50}$ which were used to ensure consistency with observational studies. For dark matter (DM) particles, we extract kinematic properties within a spherical aperture of 30 pkpc.
@@
We assume the most widely used classification to interpret galaxy kinematics (e.g. Davis et al., 2011), classifying as aligned those galaxies with stellar-gas kinematic angles $30^{\circ}$, misaligned those with angles $30^{\circ}$. Among the misaligned galaxy population, we make a further distinction between unstable misalignments with angles $30^{\circ}-150^{\circ}]$ and counter-rotation with angles $150^{\circ}$.
@@ 5.2 External properties
Following Casanueva et al. (2022), we use the halo mass, $M_{\rm{200c}}$, as a proxy to describe the large-scale environment density within which a galaxy resides (e.g. galaxy clusters with $M_{\rm{200c}}>10^{14}$ M). $M_{\rm{200c}}$, is given by the total mass bound within a radius 200 times the critical density. Among central galaxies, it follows that a larger halo mass corresponds to a larger reservoir of hot halo gas available to potentially cool onto the central galaxy (e.g. Correa et al., 2018; Mitchell & Schaye, 2022).
We define galaxy properties related to mergers using the hierarchical merger trees by Qu et al. (2017). Galaxies with more than one progenitor galaxy is indicative of a galaxy merger having occurred. We define the mass ratio of a merger as \@mathmargin
@@
where $M^{S}_{*}$} and $M^{P}_{*}$ are the stellar masses of the secondary (less massive) and primary (most massive) galaxies, respectively. Merger ratios are calculated from values in the merger tree at the point of the maximum stellar mass of the secondary galaxy within the past 2 Gyr to account for stellar stripping prior to infall (see Rodriguez-Gomez et al., 2015; Qu et al., 2017). Major and minor mergers are defined with mass ratios of $\mu_{\rm{*}}\geq 0.3$ and $0.3\geq\mu_{\rm{*}}\geq 0.1$, respectively (e.g. Cox et al., 2008; Lagos et al., 2018). We assume that ratios of $\mu_{\rm{*}}<0.1$ are direct accretion events (sometimes called micro-mergers) because these are poorly resolved in eagle (Crain et al., 2017).
@@ 5.3.1 Relaxation times
In order to estimate the relaxation time of an unstable misalignment we require galaxies to meet our sample criteria over consecutive snapshots from misalignment formation to relaxation. We also exclude unstable misalignments for which we cannot establish a formation and relaxation point within $0<z<1$. By limiting ourselves to galaxies with a complete ‘relaxation window’, we extract a smaller sample of relaxations but are able to establish a more reliable estimate of relaxation times. Additionally, our redshift range naturally places an upper limit on the relaxation time that we can accurately measure in our sample, with $t_{\mathrm{relax}}\approx 40t_{\mathrm{dyn}}$ approaching the $\approx 8$ Gyr window between $0<z<1$, for a typical ETG.
@@
For this work, we adopt a similar definition and approach as van de Voort et al. (2015), modified for use on a more diverse population of galaxies and misalignments. Our aim here is to generalise relaxation timescales of unstable misalignments for comparison with theoretical predictions (see Section 2), complementing the approach of Khim et al. (2021). As an unstable misalignment may be formed from an existing long-lived counter-rotating system, we define a separate kinematically-stable regime of $\psi_{\rm{3D}}<20^{\circ}$ and $\psi_{\rm{3D}}>160^{\circ}$. This is motivated by visual inspection of typical galaxies that are found to mostly reside within this regime, with only few outliers caused by random fluctuations. Galaxies with $20^{\circ}\leq\psi_{\rm{3D}}\leq 160^{\circ}$ are defined as kinematically-unstable. We then require any galaxy leaving the kinematically-stable regime to enter the unstable misaligned regime ($30^{\circ}\leq\psi_{\rm{3D}}\leq 150^{\circ}$) in order to be classified as significant deviation from the kinematically-stable regime [^3] Following our existing method, we note that changing the stable regime limits to $\psi_{\rm{3D}}<30^{\circ}$ and $\psi_{\rm{3D}}>150^{\circ}$, and the unstable misaligned regime to $45^{\circ}\leq\psi_{\rm{3D}}\leq 135^{\circ}$ does not meaningfully change our results.. In most cases, the first snipshot in the unstable regime also corresponds to the first snipshot in the unstable misaligned regime. We define the beginning of a relaxation from the first snipshot in the kinematically unstable regime.
As some misalignments may experience a small degree of scatter during relaxation to the stable regime, we introduce a new parameter, the latency time, $t_{\rm{lat}}$. This is defined as the minimum amount of time a galaxy must have spent in the stable regime for the unstable misalignment to have truly settled. We set this value to $t_{\rm{lat}}=0.1$ Gyr ($\gtrsim 1$ snipshots), again from visual inspection. We define the end of a relaxation as the first snipshot upon returning to the stable regime, so long as the galaxy remains in the stable regime for period of at least another $t_{\rm{lat}}$. The relaxation time, $t_{\rm{relax}}$, is then defined as the time between the snipshots corresponding to the beginning and end of the misalignment. This gives us an error of approximately $\pm 0.12$ Gyr on $t_{\rm{relax}}$.
By applying this criterion, we aim to find a balance between extracting numerous short-duration unstable misalignments that may be associated with small fluctuations from a larger overall relaxation, and extracting clearly separate unstable misalignment events as one long relaxation. We demonstrate the reliability of this criterion by showing two example galaxies in Section 6.2 and discuss its implications on our relaxation timescale distribution in Appendix A.
@@ 5.3.2 Dynamical and torquing times
The $t_{\rm{torque}}$ of a galaxy can vary between snipshots due to changes in the gas disc radius and stellar ellipticity. The gas disc radius and stellar ellipticity were found to vary within a standard deviation of $\sim 10$ per cent between misalignment formation and relaxation in our sample. We express the relaxation time in terms of time passed in Gyr and as normalised values using the the time-averaged dynamical and torquing time taken over the duration of the misalignment, denoted as $\bar{t}_{\rm{dyn}}$ and $\bar{t}_{\rm{torque}}$ , respectively. Following theoretical predictions (see Section 2), normalised relaxation timescales above $t_{\rm{relax}}/\bar{t}_{\rm{torque}}\gtrsim 1$ would be indicative of processes acting to preserve the duration of the misalignment. To account for approximations made, we define long-duration unstable misalignments as those with $t_{\rm{relax}}/\bar{t}_{\rm{torque}}>3$.
@@ 6.1 Misalignment distributions
For a brief comparison of distribution angles with observations, we make use of projected misalignment angles as viewed from the $z$-axis ($\psi_{\rm{2D}}$), within a projected half-mass radii ($r_{50,z}$) in the xyx-y plane, though our choice of axis is arbitrary as galaxies within eagle have no preferential orientation.
Figure 2 shows the projected and 3D misalignment distributions for LTGs (top) and ETGs (bottom) at $z=0.1$, with uncertainties given as Poisson errors. We include recent observational results from Bryant et al. (2019) that were obtained using the SAMI galaxy survey (Croom et al., 2012; Bryant et al., 2015).
@@
For values in projection, we find $38.9\pm 5.3$ per cent of ETGs and $7.9\pm 1.7$ per cent of LTGs display stellar-gas kinematic misalignments. For comparison, observational misalignment fractions by Bryant et al. (2019) are reported as $\approx 33$ per cent (ETGs) and $\approx 5$ per cent (LTGs), but with a stellar mass range covering a larger range of galaxy masses $10^{8}\lesssim M_{*}\lesssim 10^{12}$ M. Both the fractions of misalignments and the general shape of the misalignment distributions are in good agreement with recent observational results by Bryant et al. (2019) with a small discrepancy for LTGs within $\psi<30^{\circ}$. We find that eagle LTGs tend to be more strongly aligned within $\psi_{\rm{2D}}<10^{\circ}$, while producing fewer LTGs with $10^{\circ}<\psi_{\rm{2D}}<30^{\circ}$. This difference could be caused by a combination of the different morphological classifications used in observations, the different constraints on misalignment angle uncertainties used between our results and those of observations, and/or the different mass ranges considered. Our overall results are also in good agreement with a number of other observational studies (e.g. Davis et al., 2011; Barrera-Ballesteros et al., 2014, 2015; Chen et al., 2016; Jin et al., 2016; Ristea et al., 2022; Raimundo et al., 2023), as well as with predictions from cosmological simulations (e.g. Khim et al., 2020; Casanueva et al., 2022).
@@
We find a small peak in the counter-rotating region for both morphologies, with projected counter-rotating fractions of $10.9\pm 1.3$ per cent and $3.3\pm 0.5$ per cent for ETGs and LTGs, respectively. These were not reported in previous results in eagle (see Casanueva et al., 2022), or in the horizon-agn simulation (see Khim et al., 2020). We visually inspected our sample of $\psi_{\rm{2D}}>170^{\circ}$ misalignments and confirmed these systems to be physical and misaligned in both 3D and projection. As a whole, we find counter-rotation present in $5.7\pm 0.6$ per cent of our sample. To test the significance of this result, we perform a Kolmogorov–Smirnov (KS) test in the range $135^{\circ}-180^{\circ}]$ and compare this to a uniform distribution in the same range. We obtain KS-statistic =0.19=0.19 and p-value $1.8\times 10^{-9}$ for ETGs and KS-statistic =0.41=0.41 and p-value $3.3\times 10^{-18}$ for LTGs, indicating a statistically significant population of counter-rotators.
@@
Figure 2: Projected 2D (cyan/orange) and 3D (blue/red) stellar-gasSF misalignment angle distributions for LTGs (top) and ETGs (bottom) for our sample of misalignments at $z=0.1$. The vertical dashed black line denotes the boundary between aligned ($30^{\circ}$) and misaligned galaxy populations ($30^{\circ}$). Observational results from Bryant et al. (2019) are included in grey and should be compared against $\psi_{\rm{2D}}$ distributions. Errors are given as Poisson uncertainties. The simulated and observed galaxies exhibit fairly similar distributions.
@@
We find a 3D misalignment fraction of $44.9\pm 5.4$ per cent for ETGs and $7.2\pm 0.5$ per cent for LTGs. The overall shape of the $\psi_{\rm{3D}}$ misalignment distribution is similar to that of projected values, being almost identical for LTGs but with a smaller fraction of aligned ETGs in the range $0^{\circ}-10^{\circ}]$. We also see a small decrease in the fraction of counter-rotating systems when measured in 3D, but the number of counter-rotating systems remains significant with respect to to a uniform distribution using a KS test. These results suggest that although bias from projected viewing angles can inflate the number of aligned ETGs, the use of projected angles does not affect the overall shape or significantly reduce the number of misaligned systems.
@@
To justify our decision to consider both the co-rotating and counter-rotating regimes as stable, we show in Figure 3 how the fraction of aligned, misaligned, and counter-rotating systems varies for all snipshots between $0<z<1$. We consider the morphology independently and percentages are given with respect to the total number of galaxies of that morphological sub-sample.
@@
Figure 3: Stellar-gasSF misalignment angle fractions for LTGs (solid lines) and ETGs (dash-dotted lines) between $0<z<1$. Percentages are given with respect to the total galaxies of a given morphology at a given redshift. We classify galaxies as aligned with $\psi_{\rm{3D}}<30^{\circ}$ (blue), misaligned galaxies with $\psi_{\rm{3D}}>30^{\circ}$ (red), and counter-rotating galaxies with $\psi_{\rm{3D}}>150^{\circ}$ (purple). Errors are given as Poisson uncertainties (shaded regions). We see a slight upward trend in the misalignment fraction for both ETGs and LTGs across this redshift range, coinciding with an increase in counter-rotating fraction for both galaxy populations.
For both ETGs and LTGs we see a slight upward trend in the misalignment fraction toward $z=0$. This increase appears to result overwhelmingly from the build-up of long-lived counter-rotating systems. This is seen most clearly for ETGs, in which the $\approx 5.1$ per cent increase in misalignment fraction is captured by an $\approx 5.7$ per cent increase in counter-rotating fraction. The build-up of these systems appears to indicate that counter-rotating systems are in fact long-lasting and dynamically stable, in agreement with theoretical predictions (Tohline et al., 1982; Lake & Norman, 1983). We thus decide to treat these as a separate stable regime among the misaligned galaxy population. Some fraction of these systems may return to a co-rotating state upon another stochastic gas accretion event or undergo an apparent stellar angular momentum flip as the gasSF forms a counter-rotating stellar disc that becomes dominant. For this latter process to occur, however, the masses of the counter-rotating gas and the stars have to be similar and the system must remain counter-rotating for long enough to form a dominant population of counter-rotating stars (for an observational counterpart see e.g. NGC 4550; Rubin et al., 1992; Rix et al., 1992; Wiklind & Henkel, 2001). We note that we see these apparent stellar angular momentum flips for several of our galaxies, as shown and discussed in Section 6.2.